# Confidence Calibration in Vision-Language-Action Models

### v1

@@ Abstract
Trustworthy robot behavior requires not only high levels of task success but also that the robot can reliably quantify how likely it is to succeed. To this end, we present the first systematic study of confidence calibration in vision-language-action (VLA) foundation models, which map visual observations and natural-language instructions to low-level robot motor commands. We begin with extensive benchmarking to understand the critical relationship between task success and calibration error across multiple datasets and VLA variants, finding that task performance and calibration are not in tension. Next, we introduce prompt ensembles for VLAs, a lightweight, Bayesian-inspired algorithm that averages confidence across paraphrased instructions and consistently improves calibration. We further analyze calibration over the task time horizon, showing that confidence is often most reliable after making some progress, suggesting natural points for risk-aware intervention. Finally, we reveal differential miscalibration across action dimensions and propose action-wise Platt scaling, a method to recalibrate each action dimension independently to produce better confidence estimates. Our aim in this study is to begin to develop the tools and conceptual understanding necessary to render VLAs both highly performant and highly trustworthy via reliable uncertainty quantification.
@@ 1 Introduction
Confidence calibration, or the degree to which a model’s predicted probabilities reflect the true likelihood of its predictions being correct, is a cornerstone of reliable machine learning systems (Guo et al., 2017). When a model is well-calibrated, downstream decision-makers (e.g., humans, planning algorithms, or safety monitors) can trust that a confidence estimate of 95% implies that the predicted outcome will occur roughly 95% of the time (see Figure 1). Mismatches between confidence and actual outcomes can have severe negative consequences in high-stakes settings such as medical diagnosis or autonomous driving, motivating a rich literature on improving calibration in deep learning models (Minderer et al., 2021; Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Tian et al., 2023; Lin et al., 2024; Kadavath et al., 2022).
@@
Recently, the field of robotics has embraced a new class of vision-language-action (VLA) foundation models (Zitkovich et al., 2023; Kim et al., 2025; Black et al., 2024, 2025). Leveraging large-scale, multistage pretraining, these models translate visual observations and natural language instructions into low-level joint-space commands, unifying multimodal perception with motor control. Although still a relatively new paradigm, current VLA systems already demonstrate previously unattainable generalization across environments, tasks, and robot embodiments. Because these models encode broad semantic and visuomotor priors from large-scale pretraining, they can be efficiently fine-tuned to new robots and downstream tasks, yielding large gains over training from scratch.
@@
Contributions. To bridge this gap, we present (to our knowledge) the first systematic study of calibration in VLAs, identifying key open questions and introducing practical remedies for miscalibration. Our contributions include:
@@
    Benchmarking task success vs. calibration. We perform a comprehensive evaluation to measure the relationship between task success and calibration error across multiple datasets and VLA variants, finding that better task performance is associated with better calibration.
@@
    Action-wise scaling. We discover systematic over-/underconfidence in different action dimensions and propose a method to recalibrate each action dimension independently to produce more reliable confidence estimates.
@@ 2.1 Confidence Calibration
A model is considered well-calibrated when the confidence (i.e., probability) it assigns to an outcome matches the long-run frequency of that outcome. Deviation from this condition is often quantified via expected calibration error (ECE) (Guo et al., 2017), which is approximated by grouping predictions with similar confidence into bins and measuring the absolute difference between confidence and accuracy in each bin. Beyond ECE, other metrics such as maximum calibration error (Guo et al., 2017), Brier score (Brier, 1950), and negative log-likelihood (NLL) are used to capture related and complementary notions of miscalibration. Despite their impressive accuracy and trends toward better calibration, modern neural networks still display some persistent miscalibration (Minderer et al., 2021), especially when the task distribution does not match the training distribution (Ovadia et al., 2019). Consequently, a rich toolbox of post hoc fixes has emerged: Platt scaling (Platt, 1999), temperature scaling (Guo et al., 2017), histogram binning (Zadrozny and Elkan, 2001), and other recalibration methods (Kumar et al., 2019; Naeini et al., 2015; Zollo et al., 2024) can be applied to an already trained model to shrink residual calibration error without altering its decision rule. Other popular methods to improve calibration of neural networks include dropout (Gal and Ghahramani, 2016) and ensembling (Lakshminarayanan et al., 2017).
@@ 2.2 Vision-Language-Action Models
Vision-Language-Action models take visual data and natural language instructions as input and output robot actions (Zitkovich et al., 2023; Kim et al., 2025; Black et al., 2024, 2025). They are typically initialized from a visually-conditioned language model (VLM) that, in turn, is initialized from a pretrained large language model (LLM). This allows VLAs to leverage rich multimodal priors in joining perception and action generation into a single end-to-end pipeline. While some VLAs retain the token-based output paradigm inherited from these base models (Zitkovich et al., 2023; Kim et al., 2025), others have augmented the architecture with, e.g., flow matching action experts for smooth, high-frequency control (Black et al., 2024, 2025). These systems already perform complex, language-specified tasks across diverse environments and robot embodiments. However, they lack a reliable mechanism for quantifying the uncertainty of their chosen action sequences.
@@ 2.3 Calibration in LLMs
Given that most state-of-the-art VLAs are built on LLM backbones, it is natural to consider how our work relates to existing studies of calibration in LLMs, an area that has received considerable attention recently. Typical approaches to generating confidence estimates with LLMs include expressing calibration as a multiple-choice question for the LLM (Kadavath et al., 2022), directly verbalizing confidence with the model’s text output (Lin et al., 2022; Tian et al., 2023; Band et al., 2024), or measuring semantic consistency across many sampled outputs for the same query (Kuhn et al., 2023; Duan et al., 2024; Chen et al., 2024). However, it is difficult to adapt these methods to the VLA setting. For example, while LLMs may produce the entire sequence before measuring confidence, in robotics potential failures must be flagged much earlier in the trajectory to ensure safety and avoid costly accidents. Also, sampling multiple full trajectories in the physical world may be impossible. Finally, current VLAs lack the flexible and robust text-to-text interface of LLMs, and thus cannot be expected to, e.g., answer natural language questions about their confidence in an action prediction.
@@ 2.4 Uncertainty Quantification in Robotics
Uncertainty quantification has long been a focus in robotics, particularly through the lens of probabilistic robotics (Thrun et al., 2005; Deisenroth and Rasmussen, 2011). Recent work explores the use of model ensembles in control systems, examining how disagreements among multiple models can capture uncertainty and guide more cautious policy updates and actions (Chua et al., 2018; Pathak et al., 2019; Ataei and Dhiman, 2024). Other advances include uncertainty-based active exploration (Wang et al., 2024) and the application of conformal methods, which can provide distribution-free guarantees for planning or classification pipelines (Ren et al., 2023; Sun et al., 2023). How to obtain calibrated, task-level uncertainty estimates given the expressive, multimodal perception and control abilities of VLAs remains an open and safety-critical problem.
@@ 3 Calibration in Vision-Language-Action Models
In this section, we formalize the problem of confidence calibration for vision-language-action (VLA) policies. We define calibration, describe how contemporary VLA architectures generate actions and how to extract a corresponding confidence estimate, and introduce the standard calibration metrics used in our empirical study. Finally, we present two lightweight remedies for the miscalibration we observe in practice: prompt ensembles and action-wise Platt scaling.
@@ 3.2 Vision-Language-Action Models
At task timestep $t$, a VLA policy $\pi_{\theta}$ has access to the observed history $o_{t}=(v_{\leq t},l_{\leq t},a_{<t})$, where $v$ is a visual observation, $l$ is the (possibly fixed) natural language instruction, and $a$ is an action. The policy induces a distribution over the next action $\pi_{\theta}(a\mid o_{t}).$
When performing a task, the policy executes the most likely action $a_{t}^{\star}=\mathop{\rm argmax}_{a}\pi_{\theta}(a\mid o_{t})$. Given the need for calibration, we extract a scalar confidence score $c_{t}$ from the policy. We interpret $c_{t}$ as the policy’s estimate of the probability that the task will ultimately succeed given the observation history and the chosen action, i.e., $\mathbb{P}(Y=1\mid o_{t},a_{t}^{\star})$. Our subsequent analysis and metrics treat $c_{t}$ as a black-box number, so the following sections apply equally to token-based, diffusion-based, or other controllers, as long as they are able to emit such a scalar.
@@ 3.3 Baseline Confidence Estimation
Many state-of-the-art VLAs, including OpenVLA (Kim et al., 2025) and RT-2 (Zitkovich et al., 2023), represent an action with $D$ discrete tokens, where each represents one dimension of the robot’s action space. For each dimension $d\in\{1,\dots,D\}$ the policy outputs logits $z^{(d)}_{t}$ and probabilities
@@
Averaging plays the same role as length-normalization in LLMs, preventing confidence from being unfairly reduced for robots with many degrees of freedom. We use this as our baseline confidence estimate in our experiments.
@@ 3.4 Measuring Calibration
One popular measure of miscalibration is expected calibration error (ECE) (Guo et al., 2017; Fisch et al., 2022):
@@
Put simply, ECE measures the expected difference between confidence and accuracy over the robot’s task data distribution. The parameter $q$ is typically set to $q\in\{1,2\}$, where $ECE_{1}$ weights deviations linearly and $ECE_{2}$ penalizes larger errors.
@@
Brier score (Brier, 1950) is a classic measure of the quality of probabilistic forecasting:
@@
Brier score is an example of a proper scoring rule (Gneiting and Raftery, 2007), meaning that it is minimized only when the predicted distribution matches the true distribution for each example (as opposed to ECE, where data is grouped by confidence level). Brier score rewards both reliability (confidence matching success) and sharpness (predictions away from the population rate).
@@
Another metric used to measure calibration (Guo et al., 2017) is negative log-likelihood (NLL):
@@ 3.5 Prompt Ensembles
Similar to the VLMs from which they are derived (Zhou et al., 2024), semantically meaningless lexical differences in instructions, e.g., “pick up the coffee cup” vs. “grab the mug”, might shift visual attention, alter path planning, and thereby change the predictions and confidence scores emitted by VLA models. Although such sensitivity can pose a challenge to reliable task execution, it also creates an opportunity to employ an ensemble-based approach (Lakshminarayanan et al., 2017) to confidence estimation. Specifically, we can treat the particular wording of an instruction as a latent random variable and marginalize over it via Bayesian model averaging. Such averaging over rephrasings reduces variance in the final confidence estimate by canceling out the noise induced by word choice. Implementing this idea as an algorithm, we can then:
@@
    Generate rephrasings. An auxiliary LLM produces $r$ semantically equivalent prompts $\mathcal{L}_{alt}=\{l_{alt}^{(1)},\dots,l_{alt}^{(r)}\}$ (see Table 1 for example).
@@
Conceptually, this prompt ensemble technique can play a similar role to full model ensembles (Lakshminarayanan et al., 2017) or inference-time dropout (Gal and Ghahramani, 2016). With batched inference, this method should incur negligible latency with added inference time complexity $\mathcal{O}(1)$.
@@ 3.6 Action-Wise Scaling
A standard remedy for miscalibration in classification models is to gather a small validation set from the task distribution and perform post hoc recalibration, learning a function that maps uncalibrated confidence estimates to calibrated ones (Platt, 1999; Naeini et al., 2015; Zadrozny and Elkan, 2001; Guo et al., 2017). Post hoc recalibrators are typically simple functions of the original confidence estimate.
One popular example of a post hoc recalibration method is Platt scaling (Platt, 1999; Kumar et al., 2019). Given confidence outputs and binary outcomes $\{(c_{i},y_{i})\}_{i=1}^{N}$ on a held-out validation set, one fits an affine transform $g(c)=\sigma(\alpha c+\beta)$ that minimizes NLL,
@@
where $\sigma(x)=1/(1+e^{-x})$. At inference each $c_{t}$ is replaced with $\tilde{c}_{t}=g(c_{t})$, with the goal of aligning confidence with the probability of task success while leaving the model’s actions unchanged.
Unlike classification models that emit one predictive distribution per sample, token-based VLAs output one predictive distribution per action dimension. Those dimensions can differ dramatically, for example because gripper open/close appears in nearly 100% of demonstrations, whereas “rotate wrist” is rarer. A single global transform therefore may not be able to correct all dimensions simultaneously.
@@ 4 Experiments
Having described the problem of confidence calibration in vision-language-action models, we next conduct an extensive empirical investigation, studying the following questions:
@@
    How are task success and calibration error related? Does a higher success rate imply better calibration? (Section 4.2)
@@
    How does calibration evolve over the task time horizon? (Section 4.4)

### v2

@@ Abstract
Trustworthy robot behavior requires not only high levels of task success but also that the robot can reliably quantify how likely it is to succeed. To this end, we present a first-of-its-kind study of confidence calibration in vision-language-action (VLA) foundation models, which map visual observations and natural language instructions to low-level robot motor commands. We establish a confidence baseline for VLAs, examine how task success relates to calibration error and how calibration evolves over time, and introduce two lightweight techniques to remedy the miscalibration we observe: prompt ensembles and action-wise Platt scaling. Our aim in this study is to begin to develop the tools and conceptual understanding necessary to render VLAs both highly performant and highly trustworthy via reliable uncertainty quantification.
@@ 1 Introduction
Confidence calibration, or the degree to which a model’s predicted probabilities reflect the true likelihood of its predictions being correct, is a cornerstone of reliable machine learning systems (guo2017calibration). When a model is well-calibrated, downstream decision-makers (e.g., humans, planning algorithms, or safety monitors) can trust that a confidence estimate of 95% implies that the predicted outcome will occur roughly 95% of the time (see Figure 1). Mismatches between confidence and actual outcomes can have severe negative consequences in high-stakes settings such as medical diagnosis or autonomous driving, motivating a rich literature on improving calibration in deep learning models (minderer2021revisiting; lakshminarayanan2017simple; gal2016dropout; tian2023just; lin2023generating; kadavath2022language).
@@
Recently, the field of robotics has embraced a new class of vision-language-action (VLA) foundation models (brohan2023rt2visionlanguageactionmodelstransfer; kim2024openvlaopensourcevisionlanguageactionmodel; black2024pi0visionlanguageactionflowmodel; black2025pi05visionlanguageactionmodelopenworld). Leveraging large-scale, multistage pretraining, these models translate visual observations and natural language instructions into low-level joint-space commands, unifying multimodal perception with motor control. Although still a relatively new paradigm, current VLA systems already demonstrate previously unattainable generalization across environments, tasks, and robot embodiments. Because these models encode broad semantic and visuomotor priors from large-scale pretraining, they can be efficiently fine-tuned to new robots and downstream tasks, yielding large gains over training from scratch.
@@
Contributions. To bridge this gap, we present (to our knowledge) the first study of calibration in VLAs, identifying key open questions and introducing practical remedies for miscalibration. Our contributions include:
@@
    Benchmarking task success vs. calibration. We evaluate the relationship between task success and calibration error across multiple benchmarks and VLA variants, finding that the model architecture and training objective may play significant roles in determining this relationship.
@@
    Action-wise scaling. We discover systematic over-/under-confidence in different action dimensions and propose a method to recalibrate each action dimension independently to produce more reliable confidence estimates.
@@ 2.1 Confidence Calibration
A model is considered well-calibrated when the confidence (i.e., probability) it assigns to an outcome matches the long-run frequency of that outcome. Deviation from this condition is often quantified via expected calibration error (ECE) (guo2017calibration), which is approximated by grouping predictions with similar confidence into bins and measuring the absolute difference between confidence and accuracy in each bin. Beyond ECE, other metrics such as maximum calibration error (guo2017calibration), Brier score (brier1950verification), and negative log-likelihood (NLL) are used to capture related and complementary notions of miscalibration. Despite their impressive accuracy and trends toward better calibration, modern neural networks still display some persistent miscalibration (minderer2021revisiting), especially when the task distribution does not match the training distribution (ovadia2019trust). Consequently, a rich toolbox of post hoc fixes has emerged: Platt scaling (Platt1999), temperature scaling (guo2017calibration), histogram binning (zadrozny2001obtaining), and other recalibration methods (kumar2020verifieduncertaintycalibration; naeini2015obtaining; zollo2024improvingpredictorreliabilityselective) can be applied to an already trained model to reduce calibration error without altering its decision rule. Other popular methods to improve calibration of neural networks include dropout (gal2016dropout) and ensembling (lakshminarayanan2017simple).
@@ 2.2 Vision-Language-Action Models
Vision-Language-Action models take visual data and natural language instructions as input and output robot actions (brohan2023rt2visionlanguageactionmodelstransfer; kim2024openvlaopensourcevisionlanguageactionmodel; black2024pi0visionlanguageactionflowmodel; black2025pi05visionlanguageactionmodelopenworld). They are typically initialized from a visually-conditioned language model (VLM) that, in turn, is initialized from a pretrained large language model (LLM). This allows VLAs to leverage rich multimodal priors to integrate perception and action generation into a single end-to-end pipeline. While some VLAs retain the token-based output paradigm inherited from these base models (brohan2023rt2visionlanguageactionmodelstransfer; kim2024openvlaopensourcevisionlanguageactionmodel), others have augmented the architecture with, e.g., flow matching action experts for smooth, high-frequency control (black2024pi0visionlanguageactionflowmodel; black2025pi05visionlanguageactionmodelopenworld). These systems already perform complex, language-specified tasks across diverse environments and robot embodiments. However, they lack a reliable mechanism for quantifying the uncertainty of their chosen action sequences.
@@ 2.3 Calibration in LLMs
Given that most state-of-the-art VLAs are built on LLM backbones, it is natural to consider how our work relates to existing studies of calibration in LLMs, an area that has received considerable attention recently. Typical approaches to generating confidence estimates with LLMs include expressing calibration as a multiple-choice question for the LLM (kadavath2022language), directly verbalizing confidence with the model’s text output (lin2022teaching; tian2023just; band2024linguistic), or measuring semantic consistency across many sampled outputs for the same query (kuhn2023semantic; duan2024shifting; chen2024inside). However, it is difficult to adapt these methods to the VLA setting. For example, while LLMs may produce the entire sequence before measuring confidence, in robotics potential failures must be flagged much earlier in the trajectory to ensure safety and avoid costly accidents. Also, sampling multiple full trajectories in the physical world may be impossible. Finally, current VLAs lack the flexible and robust text-to-text interface of LLMs, and thus cannot be expected to, e.g., answer natural language questions about their confidence in an action prediction.
@@ 2.4 Uncertainty Quantification in Robotics
Uncertainty quantification has long been a focus in robotics, particularly through the lens of probabilistic robotics (thrun2005probabilistic; deisenroth2011pilco). Recent work explores the use of model ensembles in control systems, examining how disagreements among multiple models can capture uncertainty and guide more cautious policy updates and actions (chua2018deep; pathak2019selfsupervisedexplorationdisagreement; ataei2024dadeewellcalibrateduncertaintyquantification). Other advances include uncertainty-based active exploration (wang2023coplanner) and the application of conformal methods, which can provide distribution-free guarantees for planning or classification pipelines (ren2023robotsaskhelpuncertainty; sun2023conformal). How to obtain calibrated, task-level uncertainty estimates given the expressive, multimodal perception and control abilities of VLAs remains an open and safety-critical problem.
@@ 3 Calibration in Vision-Language-Action Models
In this section, we formalize the problem of confidence calibration for vision-language-action models. We define calibration, describe a typical example of how contemporary VLA architectures generate actions and how to extract a corresponding confidence estimate, and introduce the standard calibration metrics used in our empirical study (see Figure 2 for a visual summary of this material). Finally, we present two lightweight remedies for the miscalibration we observe in practice: prompt ensembles and action-wise Platt scaling.
@@ 3.2 Vision-Language-Action Models
At task timestep $t$, a VLA policy $\pi_{\theta}$ has access to the observed history $o_{t}=(v_{\leq t},l_{\leq t},a_{<t})$, where $v$ is a visual observation, $l$ is the (possibly fixed) natural language instruction, and $a$ is an action. The policy induces a distribution over the next action $\pi_{\theta}(a\mid o_{t}).$ When performing a task, the policy executes the most likely action $a_{t}^{\star}=\mathop{\rm argmax}_{a}\pi_{\theta}(a\mid o_{t})$. Given the need for calibration, we must extract a scalar confidence score $c_{t}$ from the model. We will interpret $c_{t}$ as an estimate of the probability that the task will ultimately succeed given the observation history and the chosen action, i.e., $\mathbb{P}(Y=1\mid o_{t},a_{t}^{\star})$.
Next, we propose a baseline method for extracting $c_{t}$ from a broad class of token-based VLAs. However, our subsequent analysis treats $c_{t}$ as a black-box number, so the following sections apply equally to token-based, diffusion-based, or other controllers, as long as they are able to emit such a scalar.
@@ 3.3 Baseline Confidence Estimation
Many state-of-the-art VLAs, including OpenVLA (kim2024openvlaopensourcevisionlanguageactionmodel) and RT-2 (brohan2023rt2visionlanguageactionmodelstransfer), represent an action with $D$ discrete tokens, where each represents one dimension of the robot’s action space. For each dimension $d\in\{1,\dots,D\}$ the policy outputs logits $z^{(d)}_{t}$ and probabilities
@@
Averaging plays a role similar to length-normalization in LLMs (kuhn2023semantic), preventing confidence from being unfairly reduced for robots with many degrees of freedom.
@@ 3.4 Measuring Calibration
One popular measure of miscalibration is expected calibration error (ECE) (guo2017calibration; fisch2022selective):
@@
Put simply, ECE measures the expected difference between confidence and accuracy over the robot’s task data distribution. The parameter $q$ is typically set to $q\in\{1,2\}$, where $\text{ECE}_{1}$ weights deviations linearly and $\text{ECE}_{2}$ penalizes larger errors.
@@
Brier score (brier1950verification) is a classic measure of the quality of probabilistic forecasting:
@@
Brier score is an example of a proper scoring rule (Gneiting01032007), meaning that it is minimized only when the predicted distribution matches the true distribution for each example (as opposed to ECE, where data is grouped by confidence level). Brier score rewards both reliability (confidence matching success) and sharpness (predictions away from the population rate).
@@
Another metric used to measure calibration (guo2017calibration) is negative log-likelihood (NLL):
@@ 3.5 Prompt Ensembles
Similar to the VLMs from which they are derived (zhou2024analyzingmitigatingobjecthallucination), semantically meaningless lexical differences in instructions, e.g., “pick up the coffee cup” vs. “grab the mug”, might shift visual attention, alter path planning, and thereby change the predictions and confidence scores emitted by VLA models. Although such sensitivity can pose a challenge to reliable task execution, it also creates an opportunity to employ an ensemble-based approach (lakshminarayanan2017simple) to confidence estimation. Specifically, we can treat the particular wording of an instruction as a latent random variable and marginalize over it via Bayesian model averaging. Such averaging over rephrasings reduces variance in the final confidence estimate by canceling out the noise induced by word choice. Implementing this idea as an algorithm, we can then:
@@
    Generate rephrasings. An auxiliary LLM produces $r$ semantically equivalent prompts $\mathcal{L}_{alt}=\{l_{alt}^{(1)},\dots,l_{alt}^{(r)}\}$ (see Table 1 for examples).
@@
Conceptually, this prompt ensemble technique can play a similar role to full model ensembles (lakshminarayanan2017simple) or inference-time dropout (gal2016dropout). With efficient batching on parallel hardware, this method can add little (i.e., roughly constant) wall-clock latency in practice.
@@ 3.6 Action-Wise Scaling
A standard remedy for miscalibration in classification models is to gather a small validation set from the task distribution and perform post hoc recalibration, learning a function that maps uncalibrated confidence estimates to calibrated ones (Platt1999; naeini2015obtaining; zadrozny2001obtaining; guo2017calibration). Post hoc recalibrators are typically simple functions of the original score function (e.g., logits) or confidence estimate.
One popular example of a post hoc recalibration method is Platt scaling (Platt1999; kumar2020verifieduncertaintycalibration). Given confidence outputs and binary outcomes $\{(c_{i},y_{i})\}_{i=1}^{N}$ on a held-out validation set, one fits an affine transform $g(c)=\sigma(\alpha c+\beta)$ that minimizes NLL,
@@
where $\sigma(x)=1/(1+e^{-x})$. At inference, each $c_{t}$ is replaced with $\tilde{c}_{t}=g(c_{t})$, with the goal of aligning confidence with the probability of task success while leaving the model’s actions unchanged.
Unlike classification models that emit one predictive distribution per sample, many token-based VLAs output one predictive distribution per action dimension. Those dimensions can differ dramatically, for example because gripper open/close appears in nearly 100% of demonstrations, whereas “rotate wrist” is rarer. A single global transform therefore may not be able to correct all dimensions simultaneously.
@@ 4 Experiments
Having described the problem of confidence calibration in vision-language-action models, we next conduct an empirical investigation of the following questions:
@@
    How are task success and calibration error related, and what factors might play a role in this relationship? (Section 4.2)
@@
    How does calibration evolve over the task time horizon, and what does this suggest for potential monitoring systems? (Section 4.4)